# AltLoRA: Towards Better Gradient Approximation in Low-Rank Adaptation with Alternating Projections

### v1

@@ 3 Preliminary
Here, we omit the term related to $\eta^{2}.$ Compared with the full gradient update $-\eta\nabla_{W}L$, LoRA’s gradient can approximate the full gradient as long as $sB(\nabla_{A}L)+s(\nabla_{B}L)A$ is close to $\nabla_{W}L$. With a similar motivation, some previous work analyzes the approximation based on the Frobenius norm ([61, 62, 47]). Noticeably, LoRA-Pro [62] achieves gradient approxiation by adjusting the gradients of matrices $A$ and $B$ based on the following solutions:
@@ 4.1 Alternately Approximating the Full Gradient via Low-Rank Adaptation
We propose an alternating update scheme, where we update $A$ first and then update $B$ based on the new $A$. Define the low-rank modules as $A_{t}$ and $B_{t}$ at the $t$-th iteration, and the approximated gradients as ~$\tilde{\nabla}_{A}L$ and ~$\tilde{\nabla}_{B}L$, respectively. We begin by obtaining the optimal scaling gradient of $A$ by solving
@@
~$\begin{split}\tilde{\nabla}_{A_{t}}L&=\frac{1}{s}(B_{t}^{T}B_{t})^{-1}B_{t}^{T}(\nabla_{W_{t}}L)=\frac{1}{s^{2}}(B_{t}^{T}B_{t})^{-1}\nabla_{A_{t}}L\\ \tilde{\nabla}_{B_{t}}L&=\frac{1}{s}\left(\nabla_{W_{t+\frac{1}{2}}}L\right)A_{t+1}^{T}(A_{t+1}A_{t+1}^{T})^{-1}=\frac{1}{s^{2}}\nabla_{B_{t}}L(A_{t+1}A_{t+1}^{T})^{-1},\end{split}$ (8)
@@
Theorem 1 shows that both problems admit unique optimal solutions for ~$\tilde{\nabla}_{A_{t}}L$ and ~$\tilde{\nabla}_{B_{t}}L$, which only requires full rank. Therefore, it offers a new gradient approximation with less computational cost and promotes a more efficient updating strategy. Besides, instead of accessing the full gradient like full fine-tuning, the optimal gradient approximation only requires the standard gradient of $A$ or $B$ by backpropagation at each step and calculating the inverse of a small matrix with size $r\times r$.
Theorem 1 requires that the matrix $B_{t}$ and $A_{t}$ are full rank, but in the over-parameterized cases, the assumption is hard to achieve. To alleviate it, if we penalize the Frobenius norm of these two approximated gradients, i.e., weight decay, the condition can be eliminated (see Corollary 1 in Appendix). For simplicity, in the rest of the paper, we focus on the modified gradient in (8) for analysis. The closed-form solution in (8) yields the following full model update(with gradient descent)
@@ 4.2 Proper Momentum Design within the Low-Rank Subspaces
Assume $A_{t+1}A_{t+1}^{T}$ is full-rank, i.e., rank$(A_{t+1}A_{t+1}^{T})$ = $r$. If $M_{t}^{B}$ has aligned with the low-rank space $A_{t}$ in the $t$-th iteration, by minimizing the following problem
@@
Theorem 2 shows that it’s only necessary to store two small matrices so that we can optimize momentum properly. Similar to Section 4.1, we can also remove the assumption of full rank here (see Corollary 2 in Appendix). In contrast to LoRA-Pro with full-parameter learning (Space Complexity $\mathcal{O}(kd)$), we aim to strictly satisfy the space complexity $\mathcal{O}(kr+dr)$ for parameter efficiency and keep mentum adaptively aligned with the low-rank spaces as gradient approximation does.
@@
1 for $t=0,\dots,T-1$ do
2       if $t\bmod 2=0$  then
3             Update $A$:
4             Only backpropagate w.r.t. $A_{t}$ and obtain $\nabla_{A_{t}}L$
5             ~$\tilde{\nabla}_{A_{t}}L=\frac{1}{s^{2}}(B_{t}^{\top}B_{t})^{-1}\nabla_{A_{t}}L$
6             $\tilde{M}_{t}^{A}=(B_{t}^{\top}B_{t})^{-1}B_{t}^{T}B_{t-1}M_{t-1}^{A}$
7             $M_{t}^{A}\leftarrow\beta_{1}\tilde{M}_{t}^{A}+(1-\beta_{1})\tilde{\nabla}_{A_{t}}L$
8             $A_{t+1}\leftarrow A_{t}-\eta(M_{t}^{A}+\gamma A_{t})$
9
10      else
11             Update $B$:
12             Only backpropagate w.r.t. $B_{t}$ and obtain $\nabla_{B_{t}}L$
13             ~$\tilde{\nabla}_{B_{t}}L=\frac{1}{s^{2}}\nabla_{B_{t}}L(A_{t+1}A_{t+1}^{\top})^{-1}$
14             $\tilde{M}_{t}^{B}=M_{t-1}^{B}A_{t}A_{t+1}^{\top}(A_{t+1}A_{t+1}^{\top})^{-1}$
15             $M_{t}^{B}\leftarrow\beta_{1}\tilde{M}_{t}^{B}+(1-\beta_{1})\tilde{\nabla}_{B_{t}}L$
16             $B_{t+1}\leftarrow B_{t}-\eta(M_{t}^{B}+\gamma B_{t})$
17
18
@@ 5.1 Stable Feature Learning
Assume that, with the input $x$, $BAx$ has dimension $\mathcal{O}(n)$. In Algorithm 1 or Algorithm 2, if we use the same learning rate $\eta=\mathcal{O}(1)$ to update $A$ and $B$, it would achieves stable feature learning. Moreover, without momentum in AltLoRA or AltLoRA+, the model update achieves stable feature learning as well with
@@
However, when doing joint update ([77]), the update will introduce additional across term
$\eta^{2}(\nabla_{W_{t}}L)A_{t}^{T}(A_{t}A_{t}^{T})^{-1}(B_{t}^{T}B_{t})^{-1}B_{t}^{T}(\nabla_{W_{t}}L)\in\mathcal{O}(1).$
The across term is indeed the second order term w.r.t $\eta$, but it is same magnitude as $\eta(\nabla_{W_{t}}L)Proj_{r(A_{t})}$ and $\eta Proj_{c(B_{t})}(\nabla_{W_{t}}L)$ in infinite-width NN setting.
@@ 5.2 Transformation Invariance
then $\quad\quad\quad\quad\quad\quad Proj_{c(B_{1})}=Proj_{c(B_{2})},\quad Proj_{r(A_{1})}=Proj_{r(A_{2})}$
@@ 5.3 Convergence Analysis
Following [77], we provide a convergence analysis of AltLoRA (or AltLoRA+) without momentum within the over-parameterized two-layer ReLU NN tuning problem (see Appendix D.3). We show that the convergence is independent of the condition number of the data matrix. In contrast to [77], we impose fewer assumptions to establish the convergence analysis. Notably, we don’t require the extended spectral initialization in Definition 7.3 [77]. In our experimental study, AltLoRA (AltLoRA+) can achieve superior performance with the variant of initialization used by LoRA and its variants (see Appendix E.3.2), which supports our insight empirically.
@@ 6.1 Experiments on SFT of LLM: Natural Language Generation
Training Details. We assess our methods on dialogue generation with the WizardLM dataset [68], mathematical reasoning with the MetaMathQA dataset [76], and code generation with the CodeFeedBack dataset [86] using the LLama-3.1-8B and Llama-3-8B models [15] (see Appendix E.1). We compare AltLoRA and AltLoRA+ with the pretrained model, full fine-tuning, LoRA [23], PisSSA[42], rsLoRA[29], LoRA+[19], DoRA[39], AdaLoRA[82], LoRA-GA[61], LoRA-Rite [75]and LoRA-Pro[62]. To ensure fair comparisons, we closely follow the experimental protocol established by [62]. Unless otherwise stated, we fine-tune models using default hyperparameters (if used): $\beta_{1}=0.9$, $\beta_{2}=0.999$, and zero weight decay. We adopt a cosine learning rate schedule with a warm-up ratio of 0.03. LoRA adapters are applied to $\{Q,K,V,O\}$ layers. By default, we set the rank to $r=8$ and the scaling factor to $\alpha=32$ for dialogue generation tasks, and $r=8$, $\alpha=16$ for the mathematical reasoning and code generation tasks. We carefully grid search the learning rates  [^1] See Appendix E.1 for details of learning rate grid search. We set the sequence length to 1024 and the macro batch size to 4 for math and code tasks, and macro batch size to 8 for dialogue generation. All experiments are conducted on NVIDIA A100 and NVIDIA A6000 GPUs.. To obtain a reliable estimate of model performance, we perform three runs with different random seeds and report the average and standard deviation of the results.
@@
PreTrain 5.93$\pm$0.08 51.34$\pm$1.38 36.15$\pm$1.97
Full FT 6.31$\pm$0.04 73.31$\pm$0.32 50.81$\pm$1.10
LoRA 6.06$\pm$0.02 66.11$\pm$1.43 40.31$\pm$1.34
PiSSA 5.15$\pm$0.10 67.78$\pm$1.11 42.44$\pm$1.11
rsLoRA 6.10$\pm$0.06 68.12$\pm$0.44 43.91$\pm$1.44
LoRA+ 6.40$\pm$0.06 72.33$\pm$1.33 44.10$\pm$1.38
DoRA 6.08$\pm$0.03 68.33$\pm$0.88 42.13$\pm$1.31
AdaLoRA 6.08$\pm$0.05 72.63$\pm$1.45 42.21$\pm$2.66
LoRA-GA 6.00$\pm$0.09 70.33$\pm$0.91 42.01$\pm$1.21
LoRA-Pro 6.19$\pm$0.03 73.12$\pm$0.56 43.13$\pm$1.45
LoRA-Rite 6.10$\pm$0.01 74.10$\pm$0.31 43.12$\pm$0.51
AltLoRA 6.56$\pm$0.04 74.49$\pm$0.57 45.91$\pm$1.14
AltLoRA (rank=32) 6.39$\pm$0.04 73.24$\pm$0.29 46.87$\pm$1.49
AltLoRA (rank=128) 6.27$\pm$0.01 74.11$\pm$0.21 45.41$\pm$1.65
AltLoRA+ 6.16$\pm$0.02 76.91$\pm$0.31 50.10$\pm$1.35
AltLoRA+ (rank=32) 6.10 $\pm$0.02 76.32$\pm$0.29 49.97$\pm$1.52
AltLoRA+ (rank=128) 6.07$\pm$0.03 77.08$\pm$0.83 49.77$\pm$1.58
@@ 6.2 Experiments on Natural Language Understanding
Full 86.29$\pm$0.01 93.97$\pm$0.06 80.87$\pm$0.05 93.02$\pm$0.03 86.89$\pm$0.13 88.21
LoRA 85.32$\pm$0.01 93.76$\pm$0.05 81.31$\pm$0.20 92.96$\pm$0.09 86.03$\pm$0.24 87.88
RSLoRA 85.23$\pm$0.01 93.96$\pm$0.06 81.21$\pm$0.14 93.12$\pm$0.09 86.27$\pm$0.24 87.96
DoRA 85.58$\pm$0.03 93.65$\pm$0.06 81.16$\pm$0.04 93.04$\pm$0.06 86.14$\pm$0.12 87.91
LoRA+ 85.32$\pm$0.06 93.92$\pm$0.11 81.21$\pm$0.06 92.97$\pm$0.03 86.25$\pm$0.16 87.93
PiSSA 85.87$\pm$0.04 93.84$\pm$0.06 81.90$\pm$0.05 93.16$\pm$0.09 86.64$\pm$0.12 88.28
LoRA-GA 85.70$\pm$0.09 94.11$\pm$0.18 80.57$\pm$0.20 93.18$\pm$0.06 85.29$\pm$0.24 87.77
AdaLoRA 85.45$\pm$0.11 93.92$\pm$0.09 80.31$\pm$0.05 91.66$\pm$0.05 86.16$\pm$0.60 87.50
LoRA-Pro 85.70$\pm$0.11 93.92$\pm$0.10 78.42$\pm$0.03 93.15$\pm$0.03 86.54$\pm$0.50 87.55
AltLoRA 85.26$\pm$0.04 93.87$\pm$0.05 80.44$\pm$0.09 91.56$\pm$0.01 86.60$\pm$0.99 87.55
AltLoRA+ 85.81$\pm$0.03 94.03$\pm$0.12 81.44$\pm$0.30 92.99$\pm$0.03 87.25$\pm$1.12 88.30
@@ 6.3 Ablation Study
Figure 1 presents an ablation study of the learning rate $\eta$ and the scaling factor $\alpha$ for LoRA, AltLoRA and AltLoRA+, using the LLaMA 3.1-8B model on mathematical reasoning tasks. The results show that our proposed methods are robust in learning rate and the scaling factor with consistent superior performance. Moreover, it shows that $\alpha=16$ obtains overall better performance compared to $\alpha=8$ and $\alpha=32$. The influence of increasing rank is reported in Table 2.

### v2

@@ 2 Preliminary
Here, we omit the term related to $\eta^{2}.$ Compared with the full gradient update $\eta\nabla_{W}L$, LoRA’s gradient can approximate the full gradient as long as $sB(\nabla_{A}L)+s(\nabla_{B}L)A$ is close to $\nabla_{W}L$. With a similar motivation, some previous work analyzes the approximation based on the Frobenius norm ([65, 66, 50]). Noticeably, LoRA-Pro [66] achieves gradient approxiation by adjusting the gradients of matrices $A$ and $B$ based on the following solutions:
@@ 3.1 Alternately Approximating the Full Gradient via Low-Rank Adaptation
We propose an alternating update scheme, where we update $A$ first and then update $B$ based on the new $A$. Define the low-rank modules as $A_{t}$ and $B_{t}$ at the $t$-th iteration, and the approximated gradients as $\tilde{\nabla}_{A}L$ and $\tilde{\nabla}_{B}L$, respectively. We begin by obtaining the optimal scaling gradient of $A$ by solving
@@
$\begin{split}\tilde{\nabla}_{A_{t}}L&=\frac{1}{s}(B_{t}^{T}B_{t})^{-1}B_{t}^{T}(\nabla_{W_{t}}L)=\frac{1}{s^{2}}(B_{t}^{T}B_{t})^{-1}\nabla_{A_{t}}L\\ \tilde{\nabla}_{B_{t}}L&=\frac{1}{s}\left(\nabla_{W_{t+\frac{1}{2}}}L\right)A_{t+1}^{T}(A_{t+1}A_{t+1}^{T})^{-1}=\frac{1}{s^{2}}\nabla_{B_{t}}L(A_{t+1}A_{t+1}^{T})^{-1},\end{split}$ (8)
@@
Theorem 1 shows that both problems admit unique optimal solutions for $\tilde{\nabla}_{A_{t}}L$ and $\tilde{\nabla}_{B_{t}}L$, which only requires full rank. Therefore, it offers a new gradient approximation with less computational cost and promotes a more efficient updating strategy. Besides, instead of accessing the full gradient like full fine-tuning, the optimal gradient approximation only requires the standard gradient of $A$ or $B$ by backpropagation at each step and calculating the inverse of a small matrix with size $r\times r$.
Theorem 1 requires that the matrix $B_{t}$ and $A_{t}$ are full rank, but in the over-parameterized cases, the assumption is hard to achieve. To alleviate it, if we penalize the Frobenius norm of these two approximated gradients, i.e., weight decay, the condition can be eliminated (see Corollary 1). For simplicity, in the rest of the paper, we focus on the modified gradient in (8) for analysis. The closed-form solution in (8) yields the following full model update(with gradient descent)
@@ 3.2 Proper Momentum Design within the Low-Rank Subspaces
[…]
@@
Theorem 2 shows that it’s only necessary to store two small matrices so that we can optimize momentum properly. Similar to Section 3.1, we can also remove the assumption of full rank here (see Corollary 2). In contrast to LoRA-Pro with full-parameter learning (Space Complexity $\mathcal{O}(kd)$), we aim to strictly satisfy the space complexity $\mathcal{O}(kr+dr)$ for parameter efficiency and keep mentum adaptively aligned with the low-rank spaces as gradient approximation does.
@@
1exfor $t=0,\dots,T-1$ do
 if $t\bmod 2=0$ then
    Update AA:
      Only backpropagate w.r.t. $A_{t}$ and obtain $\nabla_{A_{t}}L$
    $\tilde{\nabla}_{A_{t}}L=\frac{1}{s^{2}}(B_{t}^{\top}B_{t})^{-1}\nabla_{A_{t}}L$
    $\tilde{M}_{t}^{A}=(B_{t}^{\top}B_{t})^{-1}B_{t}^{T}B_{t-1}M_{t-1}^{A}$
    $M_{t}^{A}\leftarrow\beta_{1}\tilde{M}_{t}^{A}+(1-\beta_{1})\tilde{\nabla}_{A_{t}}L$
    $A_{t+1}\leftarrow A_{t}-\eta(M_{t}^{A}+\gamma A_{t})$
 else
    Update $B$:
      Only backpropagate w.r.t. $B_{t}$ and obtain $\nabla_{B_{t}}L$
    $\tilde{\nabla}_{B_{t}}L=\frac{1}{s^{2}}\nabla_{B_{t}}L(A_{t+1}A_{t+1}^{\top})^{-1}$
    $\tilde{M}_{t}^{B}=M_{t-1}^{B}A_{t}A_{t+1}^{\top}(A_{t+1}A_{t+1}^{\top})^{-1}$
    $M_{t}^{B}\leftarrow\beta_{1}\tilde{M}_{t}^{B}+(1-\beta_{1})\tilde{\nabla}_{B_{t}}L$
    $B_{t+1}\leftarrow B_{t}-\eta(M_{t}^{B}+\gamma B_{t})$
@@ 4.1 Stable Feature Learning
Assume that, with the input $x$, BAxBAx has dimension $\mathcal{O}(n)$. In Algorithm 1 or Algorithm 2, if we use the same learning rate $\eta=\mathcal{O}(1)$ to update $A$ and $B$, it would achieves stable feature learning. Moreover, without momentum in AltLoRA or AltLoRA+, the model update achieves stable feature learning as well with
@@
However, when doing joint update ([81]), the update will introduce additional across term $\eta^{2}(\nabla_{W_{t}}L)A_{t}^{T}(A_{t}A_{t}^{T})^{-1}(B_{t}^{T}B_{t})^{-1}B_{t}^{T}(\nabla_{W_{t}}L)\in\mathcal{O}(1)$. The across term is indeed the second order term w.r.t $\eta$, but it is same magnitude as $\eta(\nabla_{W_{t}}L)Proj_{r(A_{t})}$ and $\eta Proj_{c(B_{t})}(\nabla_{W_{t}}L)$ in infinite-width NN setting.
@@ 4.2 Transformation Invariance
then $\quad\quad\quad\quad\quad\quad Proj_{c(B_{1})}=Proj_{c(B_{2})},\quad Proj_{r(A_{1})}=Proj_{r(A_{2})}$ .
@@ 4.3 Convergence Analysis
Following [81], we provide a convergence analysis of AltLoRA (or AltLoRA+) without momentum within the over-parameterized two-layer ReLU NN tuning problem (see Appendix D.3). In Theorem 7, we show that the convergence is independent of the condition number of the data matrix. In contrast to [81], we impose fewer assumptions to establish the convergence analysis. Notably, we don’t require the extended spectral initialization in Definition 7.3 [81]. In our experimental study, AltLoRA (AltLoRA+) can achieve superior performance with the variant of initialization used by LoRA and its variants (see Appendix E.3.2), which supports our insight empirically.
@@ 5.1 Experiments on SFT of LLM: Natural Language Generation
Training Details. We assess our methods on dialogue generation with the WizardLM dataset [72], mathematical reasoning with the MetaMathQA dataset [80], and code generation with the CodeFeedBack dataset [90] using the LLama-3.1-8B and Llama-3-8B models [17] (see Appedix E.1). We compare AltLoRA and AltLoRA+ with the pretrained model, full fine-tuning, LoRA [25], PisSSA[44], rsLoRA[31], LoRA+[21], DoRA[41], AdaLoRA[86], LoRA-GA[65], LoRA-Rite [79]and LoRA-Pro[66]. To ensure fair comparisons, we closely follow the experimental protocol established by [66]. Unless otherwise stated, we fine-tune models using default hyperparameters (if used): $\beta_{1}=0.9$, $\beta_{2}=0.999$, and zero weight decay. We adopt a cosine learning rate schedule with a warm-up ratio of 0.03. LoRA adapters are applied to $\{Q,K,V,O\}$ layers. By default, we set the rank to $r=8$ and the scaling factor to $\alpha=32$ for dialogue generation tasks, and $r=8$, $\alpha=16$ for the mathematical reasoning and code generation tasks. We carefully grid search the learning rates See Appendix E.1 for details of learning rate grid search. We set the sequence length to 1024 and the macro batch size to 4 for math and code tasks, and macro batch size to 8 for dialogue generation. All experiments are conducted on NVIDIA A100 and NVIDIA A6000 GPUs.. To obtain a reliable estimate of model performance, we perform three runs with different random seeds and report the average and standard deviation of the results.
@@
[…]
@@ 5.2 Experiments on Natural Language Understanding
[…]
@@ 5.3 Ablation Study
Figure 1 presents an ablation study of the learning rate $\eta$ and the scaling factor $\alpha$ for LoRA, AltLoRA and AltLoRA+, using the LLaMA 3.1-8B model on mathematical reasoning tasks. The results show that our proposed methods are robust in learning rate and the scaling factor with consistent superior performance. Moreover, it shows that $\alpha=16$ obtains overall better performance compared to $\alpha=8$ and $\alpha=32$. The influence of increasing rank is reported in Table 2 (see Appendix E.3 of the results on Llama-3-8B model).